OpenStudy (anonymous):

What is the speed of electron in its orbital around the nucleus? Do you think the electron changes between particle-like and wave-like?

OpenStudy (anonymous):

for example the electron may have some mass. when it jumps from one orbit to another, the energy (released in terms of photons) is equal to difference between the energy levels. What if that energy released is given by E=mc^2 where the mass is given by the change in mass of the electron?

OpenStudy (unklerhaukus):

electrons are not moving in orbits around the nucleus like planets around a star, electron are found in orbitals which are probability densities electrons; like all matter, are both wavelike and particle like always, which form is observed depends on the method of observation, the mass and the velocity

OpenStudy (anonymous):

can we calculate the mass of electron while in its orbital? That is, calculating the mass while in its orbital not when it is out of the orbital?

OpenStudy (unklerhaukus):

the rest mass of an electron \[m_0 =m_e = 9.11×10^{−31} kg\] the mass of a particle moving at relativistic speeds \[m= \gamma m_0\] where \[ \gamma = {1 \over \sqrt(1-({v \over c})^2) } \] http://www.newton.dep.anl.gov/askasci/phy05/phy05222.htm

OpenStudy (anonymous):

in general the motion of an electron in an atom is non-relativistic.

OpenStudy (anonymous):

its not changing its nature. its the experiment that we are conducting , decides.

OpenStudy (anonymous):

if u believe that u r upto, then go through r. fynman,s xplanation of wave particle duality in qed.